OpenStudy (anonymous):

Which of the following is not a step used when graphing the equation y =1 over 2 x - 3 using the slope-intercept method?

hartnn (hartnn):

choices/options?

OpenStudy (anonymous):

i cant copy it but hold up

OpenStudy (anonymous):

Plot the y-intercept (0, -3). Draw a line through the two points. Count up 1 and right 2 from the y-intercept. Plot the x-intercept (6, 0)

hartnn (hartnn):

is the equation \(\huge y=\frac{1}{2x-3}\)

OpenStudy (anonymous):

noo y= 1/2x - 3

hartnn (hartnn):

ohh...now i get it.

hartnn (hartnn):

can u find the intercepts ?

OpenStudy (anonymous):

noo at all

hartnn (hartnn):

to get x-intercept, put y=0 in y= 1/2x - 3 to get y-intercept, put x=0 in y= 1/2x - 3 what u get?

OpenStudy (anonymous):

umm 3 -6

hartnn (hartnn):

y=1/2x-3, put y=0 3=1/2x x=6 x=0 in y=1/2x-3 y=-3 got this ?

OpenStudy (anonymous):

umm yea

hartnn (hartnn):

so both the intercepts are correct. also this is a straight line. so what do u think is left out from options?

OpenStudy (anonymous):

idk

hartnn (hartnn):

on.ly this is left out which is incorrect Count up 1 and right 2 from the y-intercept.